# AI-enhanced collective intelligence

### v3

@@ Introduction
In a similar vein, the healthcare sector also leverages AI-enhanced collective intelligence, addressing challenges such as diagnostic accuracy and medical data analysis, which are critical for improving patient outcomes and advancing medical research. The Human Diagnosis Project (Human Dx) 211 is an open online system of collective medical insights that provides healthcare support to patients. Human Dx enables healthcare professionals worldwide to contribute expertise and collaborate on complex diagnostic cases 212. The platform facilitates idea exchange and mutual learning to ultimately benefit patient care by utilizing machine learning techniques that automatically learn from classifying patterns in crowdsourced data. Human Dx makes the most of limited medical resources by harnessing the collective intelligence of human physicians and AI, enabling more accurate, affordable, and accessible care for those in need.
CrowdEEG 213 is a collaborative annotation tool for medical time series data. The project combines human and machine intelligence for scalable and accurate human clinical EEG data analysis. It trains machine learning algorithms using feedback from clinical experts and non-expert crowds to perform feature detection or classification tasks on medical time series data.
In healthcare, the use of AI-CI poses the risk of amplifying existing biases in human-only diagnosis once we combine AI with human judgment 214. Consequently, errors resulting from such biased decision-making can erode patients’ trust not only in the AI systems but also in the medical professionals utilizing them.
@@
AI-enhanced collective intelligence in financial services optimizes investment strategies and improve market predictions, demonstrating the powerful synergy between human expertise and machine learning. Numerai 215 is an AI-run quant hedge fund built on crowdsourced machine learning models. Numerai hosts an innovative data science tournament in which a global network of data scientists develops machine-learning models to predict stock markets. Based on the combined knowledge of the participants, Numerai combines and aggregates these crowdsourced predictive models into an ensemble model to derive investment strategies in the financial markets.
CryptoSwarm AI 216 is a forecasting service that provides rigorous insights and intelligence on cryptocurrencies and other Web3 assets. A combination of AI technology and real-time human insights powers it. CryptoSwarm AI uses Swarm AI 217 to amplify the collective intelligence of the online communities quickly, enabling networked groups to converge on optimized solutions in real time.
In financial markets, research generally indicates positive impacts on market efficiency due to the superior information processing and optimization capabilities of machines, coupled with human anticipation of these effects. However, this improvement is sometimes disrupted by rapid price spikes and crashes, resulting from the herding behavior of machines and the challenges humans face in intervening at extremely fast timescales 218.
@@
The supply chain and real estate sectors also benefit from AI-enhanced collective intelligence, improving efficiency, transparency, and decision-making through real-time data integration and analysis. MarineTraffic 219 is a web-based platform that creates a global network of vessel tracking information, enabling users to track vessels, monitor maritime traffic, analyze trends, and make informed decisions. It relies on the collective contributions of the maritime community, who voluntarily transmit real-time information about vessel positions and movements from various sources, including AIS (automatic identification system) data and satellite data. MarineTraffic employs neural network architectures for automatic maritime object detection using satellite imagery 220. More examples of AI applications are available at MarineTraffic Research 221.
Waze 222 is a GPS navigation software that works on smartphones and other computers. It analyzes crowdsourced GPS data and user-submitted information along the route in real-time and uses community editing to ensure the accuracy of the map data. Waze uses AI algorithms to predict traffic, optimize routes 223, and provide personalized recommendations by analyzing driver behaviors. Waze has implemented integration with Google Assistant, enabling users to use the application by utilizing the voice command “Hey Google” 224.
By improving efficiency, reducing waste, and enabling more equitable distribution and accessibility of resources and spaces, AI can play a significant role in addressing inequalities in access to resources, paving the way for sustainable and efficient infrastructure development 225.
@@
Finally, the agriculture sector leverages AI-enhanced collective intelligence to improve farming practices, resource management, and crop yields, addressing global food security challenges. WeFarm 226 is a social networking platform connecting the small-scale farming community. It helps millions of African farmers meet, exchange solutions to their questions, and trade equipment and supplies. Farmers ask each other questions about agriculture and promptly receive content and ideas from other farmers worldwide through crowdsourcing. Natural language processing processes and understands the text-based messages farmers send in different regions. Machine learning matching algorithms consider farmers’ needs and expertise, identify the most suitable responses, and present them to the farmer who posed the question.
Mercy Corps AgriFin 227 harnesses the power of AI and a global network of partners to transform agriculture for smallholder farmers. By utilizing state-of-the-art imagery, modeling, and analysis, AI enables precise management of crops and adaptation strategies against climate change impacts. AgriFin’s integration of digital technology and data, supported by a global network, empowers these farmers in a digitally interconnected world.
In light of the global food crisis and the quest for sustainable solutions, the importance of AI-enhanced collective intelligence platforms becomes increasingly critical. These platforms can significantly contribute to addressing agricultural challenges and transitioning towards global sustainable food systems 228.
@@
While AI can help mitigate certain human biases, they also have the possibility to introduce new ones. These biases can stem from various stages of AI development and deployment. AI systems learn from the data they are trained on, and biased training data can cause the AI to replicate these biases. Additionally, algorithmic design choices may inadvertently favor certain groups over others. Human-in-the-loop bias is another concern, where human inputs during training and feedback, such as biased labeling by annotators can be learned by the AI. Moreover, user interactions with AI systems can also influence and reinforce these biases. AI biases pose risks to fair and equitable outcomes in society, especially in sensitive areas such as healthcare and education. Bias in AI algorithms can lead to unfair treatment or misdiagnosis in healthcare, and biased educational tools can perpetuate existing inequalities. Understanding how AI biases and human biases interact, and how to avoid doubly biased decisions made by human-machine intelligence 229, are crucial areas for further study.
@@
Many AI models, particularly deep learning algorithms, operate as “black boxes” 230, making understanding how they arrive at specific decisions or recommendations difficult. This opacity can hinder trust and acceptance among users, especially in high-stakes domains like healthcare and law. Explainability is essential for ensuring accountability, enabling users to understand and challenge AI decisions, and identifying and correcting biases within the system. Developing AI models that are both effective and interpretable requires ongoing research and innovation.
@@
The application areas of AI-enhanced collective intelligence continually expand as technology advances and new opportunities emerge. As introduced in the previous section, successful examples of AI-enhanced collective intelligence exist in various domains. However, challenges coexist with opportunities. When realizing human-AI collective intelligence, many aspects must be considered, such as human-machine communication, trust, crowd retention, technology design, and ethical issues 231. Here, we briefly discuss some of these challenges and call for further work on each aspect.
@@
Clear and effective communication between humans and AI is the engine for collective intelligence to emerge 168. Communication refers to the process of exchanging information between teammates 232. It is essential for team performance as it contributes to the development and maintenance of shared mental models and the successful execution of necessary team processes 232, 233.
For effective communication in human-AI teamwork, AI must first be able to model human information comprehension accurately and then effectively communicate in a manner understandable to humans. Current AI technologies face limitations in verbal and contextual understanding 233 and cognitive capabilities, such as reasoning about others’ mental states and intentions 234. Detection, interpretation, and reasoning about social cues from a human perspective is imperative to ensure effective coordination, but AI has yet to achieve this. Although recent developments in LLMs appear promising, further advancing our understanding of how LLMs can be utilized in such a way is crucial 235, 236.
AI’s inability to interpret nonverbal cues and limited self-explanation hinder human collaboration. Addressing these challenges, research suggests the development of AI agents with cognitive architectures that can facilitate both a Machine Theory of the Human Mind (MToHM) and a Human Theory of the Machine Mind (HToMM) will be especially important for supporting the emergence of collective intelligence 168.
@@
In team success, a human member’s trust in a machine is pivotal 237. Research indicates that users interact differently with AI than humans, often exhibiting more openness and self-disclosure with human partners 238. Developing explanatory methods that foster appropriate trust in AI is challenging yet crucial for enhancing performance in human-AI teams 178. Meanwhile, over-reliance on AI, especially when faced with incorrect AI advice, can be detrimental and lead to human skill degradation. Hence, it is essential for humans to judiciously rely on AI when appropriate and exercise self-reliance in the face of inaccurate AI guidance 239, especially in high-stakes situations. More tech literacy, public understanding of AI, and advancements in the technology itself will help in overcoming these challenges.
@@
Albeit beneficial in terms of efficiency, employing AI and citizen scientists together has been reported to damage the retention of citizen scientists 240, 206. There is a trade-off between efficiency and retention of the crowds when deploying AI in citizen science projects, and ultimately, if all the crowds leave the projects, the performance of AI alone would decline again. One of the factors contributing to the retention of volunteers is reported to be the social ties they make in their teams 241. The challenge of forming social bonds with AI could increase human feelings of loneliness and isolation, potentially affecting mental health. Questions about identity and recognition could arise when humans are rewarded in conjunction with AI. Studies highlight concerns regarding the anthropomorphic appearance of social robots potentially threatening human distinctiveness 242. Additionally, collaboration with AI may demotivate humans due to a lack of competitive drive and an over-reliance on AI, which could diminish human participation and initiative. Understanding how to balance AI collaboration with maintaining human motivation and retention is crucial.
@@
The interface design of AI technology influences user experiences, engagement, and acceptance of AI-driven communication. Interfaces should be user-friendly, promoting seamless information exchange and aligning with human cognitive processes. Since AI systems need time to process human input and generate output, there may be a significant time lag between the interactions in some instances. The delay of feedback between human input and AI responses can adversely affect 243 the coordination and interaction efficiency between humans and AI, especially in collaborative tasks. On the other hand, human behavior possesses specific temporal patterns, mainly determined by circadian rhythm 244. It is essential to design such hybrid systems so that the temporal characteristics of both entities match, support mutual understanding, and balance the cognitive load and workload distribution. Moreover, these systems should be designed to augment human contributions 245. To effectively cater to diverse user groups, human-centered AI design approaches, including the use of personas 246, can be employed to develop adaptive and inclusive human-AI interfaces.
@@
When integrating AI into human interactions, several ethical considerations emerge. Addressing these concerns involves balancing benefits and minimizing potential harms. Key issues include the ethical implications of mandatory AI use, especially when humans are reluctant or do not require AI assistance. When it comes to AI transparency, a study reported that humans who were paired with bots to play a repeated prisoner’s dilemma game outperformed human-human teams initially but later started to defect once they became aware that they were playing with a bot, showing a trade-off between efficiency and transparency 247. Another critical aspect is addressing biases inherent in AI outputs, which are influenced by the data inputs. It is essential to consider how AI integration might exacerbate or introduce new biases in decision-making 229. Furthermore, liability questions arise when AI errors cause serious harm, an issue compounded by the lack of comprehensive regulations and laws. For instance, a robot failed in object recognition and crushed a human worker to death in a factory in South Korea 248. This highlights the urgent need for comprehensive legal frameworks to address such liabilities and ensure safety in environments where AI interacts with humans. As AI becomes more prevalent in the workplace, constructing ethical human-AI teams becomes a complex, evolving challenge 249.
@@ Conclusion
In current real-world AI-CI applications, AI is predominantly utilized as a technical tool to facilitate data analysis through techniques such as natural language processing and computer vision. Implementing AI not only scales up tasks but, in combination with human intelligence, leads to superior performance. Although AI is critical in some decision-making processes, humans make the ultimate decisions and contributions. It is essential to recognize that AI’s function is to support and enhance human collaborative processes rather than to replace human intelligence. Nevertheless, it is also worth considering collective intelligence emerging from a group of AI entities in future research. Past work has demonstrated that the collective behavior of autonomous machines can lead to the emergence of complex social behavior 250.
@@
Our theoretical model of human-AI collective intelligence, a multilayer network with cognition, physical, and information layers, offers valuable insights into complex phenomena in such hybrid systems. For example, in the case of the Flash Crash 2012 251, the cognition layer includes high-frequency traders (algorithms) responding to market signals, the physical layer involves the actual financial transactions, and the information layer encompasses the data streams and communication networks connecting these elements. By understanding the interplay and feedback loops within these layers, our model helps identify vulnerabilities, such as how a single large sell order propagated through the network, causing a cascade of automated responses that led to the crash 218.
Similarly, regarding the reported issue of human user retention while AI is being deployed in citizen science projects 206, our model elucidates the interaction between human volunteers and AI classifiers within the different layers. It highlights the importance of maintaining a balance between efficiency (enhanced by AI) and volunteer engagement (rooted in human cognition). This approach prepares us to preemptively address potential pitfalls, ensuring more resilient and effective collective intelligence systems in the future.
@@
Various behavioral aspects, including human behavior, AI behavior, and the interactions between humans and AI, must be studied. To better understand and harness human-AI collectives for addressing societal challenges, researchers suggest building the foundations of Social AI 48 by integrating insights from Complex Systems, Network Science, and AI.
@@
Adaptive systems can learn and adapt over time. Researchers propose studying collective adaptation 64, focusing on how human collectives adjust their cognitive strategies and social networks in response to changing problems. Extending this research to include both human and AI collective adaptation could be valuable, requiring collaboration from cognitive science, social psychology, and other related fields.
@@
Our source data and code are available on GitHub: https://github.com/cuihaosabrina/AI-enhanced-collective-intelligence.

### v4

@@ Introduction
In a similar vein, the healthcare sector also leverages AI-enhanced collective intelligence, addressing challenges such as diagnostic accuracy and medical data analysis, which are critical for improving patient outcomes and advancing medical research. The Human Diagnosis Project (humandx.org) is an open online system of collective medical insights that provides healthcare support to patients. Human Dx enables healthcare professionals worldwide to contribute expertise and collaborate on complex diagnostic cases. The platform facilitates idea exchange and mutual learning to ultimately benefit patient care by utilizing machine learning techniques that automatically learn from classifying patterns in crowdsourced data. Human Dx makes the most of limited medical resources by harnessing the collective intelligence of human physicians and AI, enabling more accurate, affordable, and accessible care for those in need.
CrowdEEG 193 is a collaborative annotation tool for medical time series data. The project combines human and machine intelligence for scalable and accurate human clinical EEG data analysis. It trains machine learning algorithms using feedback from clinical experts and non-expert crowds to perform feature detection or classification tasks on medical time series data.
In healthcare, the use of AI-CI poses the risk of amplifying existing biases in human-only diagnosis once we combine AI with human judgment 194. Consequently, errors resulting from such biased decision-making can erode patients’ trust not only in the AI systems but also in the medical professionals utilizing them.
@@
AI-enhanced collective intelligence in financial services optimizes investment strategies and improve market predictions, demonstrating the powerful synergy between human expertise and machine learning. Numerai (numerai.fund) is an AI-run quant hedge fund built on crowdsourced machine learning models. Numerai hosts an innovative data science tournament in which a global network of data scientists develops machine-learning models to predict stock markets. Based on the combined knowledge of the participants, Numerai combines and aggregates these crowdsourced predictive models into an ensemble model to derive investment strategies in the financial markets.
CryptoSwarm AI is a forecasting service that provides rigorous insights and intelligence on cryptocurrencies and other Web3 assets. A combination of AI technology and real-time human insights powers it. CryptoSwarm AI uses Swarm AI to amplify the collective intelligence of the online communities quickly, enabling networked groups to converge on optimized solutions in real time.
In financial markets, research generally indicates positive impacts on market efficiency due to the superior information processing and optimization capabilities of machines, coupled with human anticipation of these effects. However, this improvement is sometimes disrupted by rapid price spikes and crashes, resulting from the herding behavior of machines and the challenges humans face in intervening at extremely fast timescales 195.
@@
The supply chain and real estate sectors also benefit from AI-enhanced collective intelligence, improving efficiency, transparency, and decision-making through real-time data integration and analysis. MarineTraffic (marinetraffic.com) is a web-based platform that creates a global network of vessel tracking information, enabling users to track vessels, monitor maritime traffic, analyze trends, and make informed decisions. It relies on the collective contributions of the maritime community, who voluntarily transmit real-time information about vessel positions and movements from various sources, including AIS (automatic identification system) data and satellite data. MarineTraffic employs neural network architectures for automatic maritime object detection using satellite imagery 196.
Waze (www.waze.com) is a GPS navigation software that works on smartphones and other computers. It analyzes crowdsourced GPS data and user-submitted information along the route in real-time and uses community editing to ensure the accuracy of the map data. Waze uses AI algorithms to predict traffic, optimize routes 197, and provide personalized recommendations by analyzing driver behaviors. Waze has implemented integration with Google Assistant, enabling users to use the application by utilizing the voice command “Hey Google”.
By improving efficiency, reducing waste, and enabling more equitable distribution and accessibility of resources and spaces, AI can play a significant role in addressing inequalities in access to resources, paving the way for sustainable and efficient infrastructure development 198.
@@
Finally, the agriculture sector leverages AI-enhanced collective intelligence to improve farming practices, resource management, and crop yields, addressing global food security challenges. WeFarm (wefarm.co) is a social networking platform connecting the small-scale farming community. It helps millions of African farmers meet, exchange solutions to their questions, and trade equipment and supplies. Farmers ask each other questions about agriculture and promptly receive content and ideas from other farmers worldwide through crowdsourcing. Natural language processing processes and understands the text-based messages farmers send in different regions. Machine learning matching algorithms consider farmers’ needs and expertise, identify the most suitable responses, and present them to the farmer who posed the question.
Mercy Corps AgriFin (www.mercycorpsagrifin.org) harnesses the power of AI and a global network of partners to transform agriculture for smallholder farmers. By utilizing state-of-the-art imagery, modeling, and analysis, AI enables precise management of crops and adaptation strategies against climate change impacts. AgriFin’s integration of digital technology and data, supported by a global network, empowers these farmers in a digitally interconnected world.
In light of the global food crisis and the quest for sustainable solutions, the importance of AI-enhanced collective intelligence platforms becomes increasingly critical. These platforms can significantly contribute to addressing agricultural challenges and transitioning towards global sustainable food systems 199.
@@
While AI can help mitigate certain human biases, they also have the possibility to introduce new ones. These biases can stem from various stages of AI development and deployment. AI systems learn from the data they are trained on, and biased training data can cause the AI to replicate these biases. Additionally, algorithmic design choices may inadvertently favor certain groups over others. Human-in-the-loop bias is another concern, where human inputs during training and feedback, such as biased labeling by annotators can be learned by the AI. Moreover, user interactions with AI systems can also influence and reinforce these biases. AI biases pose risks to fair and equitable outcomes in society, especially in sensitive areas such as healthcare and education. Bias in AI algorithms can lead to unfair treatment or misdiagnosis in healthcare, and biased educational tools can perpetuate existing inequalities. Understanding how AI biases and human biases interact, and how to avoid doubly biased decisions made by human-machine intelligence 200, are crucial areas for further study.
@@
Many AI models, particularly deep learning algorithms, operate as “black boxes” 201, making understanding how they arrive at specific decisions or recommendations difficult. This opacity can hinder trust and acceptance among users, especially in high-stakes domains like healthcare and law. Explainability is essential for ensuring accountability, enabling users to understand and challenge AI decisions, and identifying and correcting biases within the system. Developing AI models that are both effective and interpretable requires ongoing research and innovation.
@@
The application areas of AI-enhanced collective intelligence continually expand as technology advances and new opportunities emerge. As introduced in the previous section, successful examples of AI-enhanced collective intelligence exist in various domains. However, challenges coexist with opportunities. When realizing human-AI collective intelligence, many aspects must be considered, such as human-machine communication, trust, crowd retention, technology design, and ethical issues 202. Here, we briefly discuss some of these challenges and call for further work on each aspect.
@@
Clear and effective communication between humans and AI is the engine for collective intelligence to emerge 164. Communication refers to the process of exchanging information between teammates 203. It is essential for team performance as it contributes to the development and maintenance of shared mental models and the successful execution of necessary team processes 203, 204.
For effective communication in human-AI teamwork, AI must first be able to model human information comprehension accurately and then effectively communicate in a manner understandable to humans. Current AI technologies face limitations in verbal and contextual understanding 204 and cognitive capabilities, such as reasoning about others’ mental states and intentions 205. Detection, interpretation, and reasoning about social cues from a human perspective is imperative to ensure effective coordination, but AI has yet to achieve this. Although recent developments in LLMs appear promising, further advancing our understanding of how LLMs can be utilized in such a way is crucial 206, 207.
AI’s inability to interpret nonverbal cues and limited self-explanation hinder human collaboration. Addressing these challenges, research suggests the development of AI agents with cognitive architectures that can facilitate both a Machine Theory of the Human Mind (MToHM) and a Human Theory of the Machine Mind (HToMM) will be especially important for supporting the emergence of collective intelligence 164.
@@
In team success, a human member’s trust in a machine is pivotal 208. Research indicates that users interact differently with AI than humans, often exhibiting more openness and self-disclosure with human partners 209. Developing explanatory methods that foster appropriate trust in AI is challenging yet crucial for enhancing performance in human-AI teams 174. Meanwhile, over-reliance on AI, especially when faced with incorrect AI advice, can be detrimental and lead to human skill degradation. Hence, it is essential for humans to judiciously rely on AI when appropriate and exercise self-reliance in the face of inaccurate AI guidance 210, especially in high-stakes situations. More tech literacy, public understanding of AI, and advancements in the technology itself will help in overcoming these challenges.
@@
Albeit beneficial in terms of efficiency, employing AI and citizen scientists together has been reported to damage the retention of citizen scientists 190. There is a trade-off between efficiency and retention of the crowds when deploying AI in citizen science projects, and ultimately, if all the crowds leave the projects, the performance of AI alone would decline again. One of the factors contributing to the retention of volunteers is reported to be the social ties they make in their teams 211. The challenge of forming social bonds with AI could increase human feelings of loneliness and isolation, potentially affecting mental health. Questions about identity and recognition could arise when humans are rewarded in conjunction with AI. Studies highlight concerns regarding the anthropomorphic appearance of social robots potentially threatening human distinctiveness 212. Additionally, collaboration with AI may demotivate humans due to a lack of competitive drive and an over-reliance on AI, which could diminish human participation and initiative. Understanding how to balance AI collaboration with maintaining human motivation and retention is crucial.
@@
The interface design of AI technology influences user experiences, engagement, and acceptance of AI-driven communication. Interfaces should be user-friendly, promoting seamless information exchange and aligning with human cognitive processes. Since AI systems need time to process human input and generate output, there may be a significant time lag between the interactions in some instances. The delay of feedback between human input and AI responses can adversely affect 213 the coordination and interaction efficiency between humans and AI, especially in collaborative tasks. On the other hand, human behavior possesses specific temporal patterns, mainly determined by circadian rhythm 214. It is essential to design such hybrid systems so that the temporal characteristics of both entities match, support mutual understanding, and balance the cognitive load and workload distribution. Moreover, these systems should be designed to augment human contributions 215. To effectively cater to diverse user groups, human-centered AI design approaches, including the use of personas 216, can be employed to develop adaptive and inclusive human-AI interfaces.
@@
When integrating AI into human interactions, several ethical considerations emerge. Addressing these concerns involves balancing benefits and minimizing potential harms. Key issues include the ethical implications of mandatory AI use, especially when humans are reluctant or do not require AI assistance. When it comes to AI transparency, a study reported that humans who were paired with bots to play a repeated prisoner’s dilemma game outperformed human-human teams initially but later started to defect once they became aware that they were playing with a bot, showing a trade-off between efficiency and transparency 217. Another critical aspect is addressing biases inherent in AI outputs, which are influenced by the data inputs. It is essential to consider how AI integration might exacerbate or introduce new biases in decision-making 200. Furthermore, liability questions arise when AI errors cause serious harm, an issue compounded by the lack of comprehensive regulations and laws. For instance, a robot failed in object recognition and crushed a human worker to death in a factory in South Korea 218. This highlights the urgent need for comprehensive legal frameworks to address such liabilities and ensure safety in environments where AI interacts with humans. As AI becomes more prevalent in the workplace, constructing ethical human-AI teams becomes a complex, evolving challenge 219.
@@ Conclusion
In current real-world AI-CI applications, AI is predominantly utilized as a technical tool to facilitate data analysis through techniques such as natural language processing and computer vision. Implementing AI not only scales up tasks but, in combination with human intelligence, leads to superior performance. Although AI is critical in some decision-making processes, humans make the ultimate decisions and contributions. It is essential to recognize that AI’s function is to support and enhance human collaborative processes rather than to replace human intelligence. Nevertheless, it is also worth considering collective intelligence emerging from a group of AI entities in future research. Past work has demonstrated that the collective behavior of autonomous machines can lead to the emergence of complex social behavior 220.
@@
Our theoretical model of human-AI collective intelligence, a multilayer network with cognition, physical, and information layers, offers valuable insights into complex phenomena in such hybrid systems. For example, in the case of the Flash Crash 2012 221, the cognition layer includes high-frequency traders (algorithms) responding to market signals, the physical layer involves the actual financial transactions, and the information layer encompasses the data streams and communication networks connecting these elements. By understanding the interplay and feedback loops within these layers, our model helps identify vulnerabilities, such as how a single large sell order propagated through the network, causing a cascade of automated responses that led to the crash 195.
Similarly, regarding the reported issue of human user retention while AI is being deployed in citizen science projects 190, our model elucidates the interaction between human volunteers and AI classifiers within the different layers. It highlights the importance of maintaining a balance between efficiency (enhanced by AI) and volunteer engagement (rooted in human cognition). This approach prepares us to preemptively address potential pitfalls, ensuring more resilient and effective collective intelligence systems in the future.
@@
Various behavioral aspects, including human behavior, AI behavior, and the interactions between humans and AI, must be studied. To better understand and harness human-AI collectives for addressing societal challenges, researchers suggest building the foundations of Social AI 47 by integrating insights from Complex Systems, Network Science, and AI.
@@
Adaptive systems can learn and adapt over time. Researchers propose studying collective adaptation 63, focusing on how human collectives adjust their cognitive strategies and social networks in response to changing problems. Extending this research to include both human and AI collective adaptation could be valuable, requiring collaboration from cognitive science, social psychology, and other related fields.
@@
Our source data and code are available publicly183.